OpenStudy (anonymous):

please write 3x-4y=24

OpenStudy (anonymous):

what do you mean by write

OpenStudy (anonymous):

3x-4y=24 Since 3x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3x from both sides. -4y=-3x+24 Divide each term in the equation by -4. -(4y)/(-4)=-(3x)/(-4)+(24)/(-4) Simplify the left-hand side of the equation by canceling the common factors. y=-(3x)/(-4)+(24)/(-4) Simplify the right-hand side of the equation by simplifying each term. y=(3x)/(4)-6

OpenStudy (anonymous):

3x-4y=24 Since -4y does not contain the variable to solve for, move it to the right-hand side of the equation by adding 4y to both sides. 3x=4y+24 Divide each term in the equation by 3. (3x)/(3)=(4y)/(3)+(24)/(3) Simplify the left-hand side of the equation by canceling the common factors. x=(4y)/(3)+(24)/(3) Simplify the right-hand side of the equation by simplifying each term. x=(4(y+6))/(3)

OpenStudy (anonymous):

That should cover it